Susan has an assignment in Math and its probability is 5/14. On the other hand, her Science assignment's probability is 3/7.
If the probability of assignments both in Math and Science is 1/7, what is the probability that she has EITHER an assignment in Math OR an assignment in Science?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

P(a OR b)=P(a)+P(b)-P(a AND b)
=(5/14)+(6/14)-(2/14)
=9/14

Answer by ikleyn(52797)

.

As this "problem" was worded,  it  WAS  TERRIBLE  to read it . . .